OpenStudy (anonymous):

Im having trouble understanding negative integers... could you help me? ex: (-2)^-3 =?

OpenStudy (anonymous):

(-2)^-3 = 1/((-2)^3) = -1/8

OpenStudy (savvy):

a -ve in the power means you need to take the reciprocal of the term in the base...CoCoTsoi has already showed it..!!!